Solve and check each of the following for X:
[5/(x-2)]+[6/(x+2)]=2

Answer by edjones(8007)

[5/(x-2)]+[6/(x+2)]=2
5(x+2)+6(x-2)=2(x^2-4) Multiply each side by (x+2)(x-2) to eliminate fractions.
5x+10+6x-12=2x^2-8
11x-2=2x^2-8
2x^2-11x-6=0
2x^2-12x+x-6=0
2x(x-6)+(x-6)=0 Factoring by grouping.
(2x+1)(x-6)=0
2x=-1
x=-1/2
x=6
.
Ed